Calculate the total amount of energy required to raise the temperature of 30g of ice from -8°C to 75°C.
Specific heat capacity for Ice = 2.1J/g°C
Latent heat of fusion of Ice = 335J/g
Specific heat capacity for Water = 4.2J/g°C
1
Expert's answer
2020-11-05T10:32:50-0500

"Q_1 = 8\times 30 \times 2.1 = 504 \;J" (from -8 °C to 0 °C)

"Q_2 = 75 \times 30 \times 4.2 = 9450 \;J" (from 0 °C to 75 °C)

"Q_3 = 30 \times 335 = 10050 \;J" (heat of melting)

Total heat "Q_T = 504 + 9450 + 10050 = 20004 \;J"

Answer: 20004 J